OpenStudy (anonymous):

Find a recursive rule for the sequence 2, 4, 6, 10, …

myininaya (myininaya):

hint: 2+4=? 4+6=?

myininaya (myininaya):

can you find those sums?

OpenStudy (anonymous):

the thing i dont understand is how the formula looks. im not very good at math. and ive never learned how to do recursive formulas.

myininaya (myininaya):

well can you find the sums mentioned above?

myininaya (myininaya):

2+4=? 4+6=?

myininaya (myininaya):

do you notice that you find a term by adding it's previous two terms for this sequence ?

myininaya (myininaya):

\[a_1,a_2,a_3,a_4,\cdots \\ \text{ we seen the following if you found the sums I asked you to above } \\ a_1+a_2=a_3 \\ a_2+a_3=a_4\]

myininaya (myininaya):

\[\text{ so say we want to find } a_n \text{ and we know the previous two terms are } \\ a_{n-1} \text{ and } a_{n-2}\] \[\text{ could you write what } a_n \text{ would be }\]